Systems Staff Meeting Notes
August 13, 2007
August 13, 2007
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Stephen WilliamsAnnouncements
Ongoing Systems Updates
- Tom Minton is getting a motorcycle learner's permit.
- Amy is going on a women's retreat. They are going to learn to shoot stuff.
- Tom Minton has a marksman's badge.
- Bobbie's stepdaughter, Brittany, graduated from UF this weekend. Her major was Event Management.
- Carol Turner's last day will be Sept. 28. Shelley Arlen will act as the Interim Director.
- Convocation is Wed. August 15.
- Physical Plant will be converting East from CCure to Lenel security in the fall.
- Hurricane prep - there is a link about it on the Library web page.
- Systems calendar entries - be sure to include your name when scheduling time out.
- Special orders - items we buy and give out to other departments should be special orders, not out of the Systems supply budget.
- Inventory - no improvement with the information provided.
- Training - end of the month.
- Verify Patron - running Batch File using Deep Freeze - new workstations replaced old ones that were the console. It should all tie together with a customization code. The fix is to redo clients and consoles. Amy is exploring getting it to run before school starts. Need to contact Deep Freeze support.
Software Projects Status
- East - network upgrade - we now have gigabit feed in the south tower. No major issues. Port 44 is the VLAN port for East Video. CNS will monitor it. North tower upgrade scheduled for this weekend.
- Public computer replacement - most places have been done, with the exception of a few minor ones. Need to replace GX240's with 260's.
- Memory upgrade request for Classroom 212.
- Education classroom is done. Amy has the documentation disks.
- SAN - Logan to start on it soon. The books we ordered just came in - that, interlinked with the VM Ware training - may have it configured initially by the end of the month.
- SMS/DHCP - per CNS: no problem with us running DHCP server. They will configure the port for us with forward. It would be completely internal to us. We will speak to CNS more about the details.
- Updates - 150 workstations need updates. Logan has not approved any lately (training last week). Some of them are outdated and just need to be deleted.
- Grover is doing alright - there was a performance issue on Friday - it was way too slow. We think we found a workaround.
- Tracking database - ongoing.
- Nothing new with Greenstone.
- We think the ILL switch is finished.
- National Library barcode prefix - requested approval from Lori, have not heard back.
- Systems blog - Stephen would really like it to be made live. Much discussion ensues. Will is going to run it by the Directors for approval.
Students and Department
- Endeca switch next Monday - Tom has changed the URL's in the template. He can't change the link in the content of pages. People will have to do that individually.
- Server side includes - Logan to set up a virtual server to test. Tom says okay - as long as it does not mess with the real one.
Systems Security Management
- Our new student, Robert, has started working. Please be vigilant in updating PAULA. Tom Minton notes he needs PAULA installed on his new machine.
- Meeting held.